OpenStudy (anonymous):

suppose you start a business. You buy a freezer for $200. It costs you $.35 to make a single-scoop ice cream cone. You sell each cone for $1.20. Write and solve a system of equations to find the break-even point for your business. @mathstudent55

OpenStudy (mathstudent55):

You need to find the expenses and the amount coming in and set them equal. Let's use x as the number of cones needed to break even. E = Expenses: You have a fixed expense of $200. Then you have an expense of $0.35 per cone for x cones. R = Revenue The amount coming in is $1.20 per cone for x cones.

OpenStudy (anonymous):

ok

OpenStudy (mathstudent55):

How much do the cones cost you if they are $0.35 each and you buy x number of them?

OpenStudy (anonymous):

$0.35x

OpenStudy (mathstudent55):

Great. The total expense, E is: E = 200 + 0.35x Now we need the total revenue. How much are x cones at $1.20 each worth?

OpenStudy (anonymous):

200+$1.20x

OpenStudy (mathstudent55):

No. There is no $200 fixed amount associated with the money coming in. The money coming in is only from the sales of x cones at $1.20 each, so the money coming in is simply $1.20x.

OpenStudy (anonymous):

ok

OpenStudy (mathstudent55):

Total expenses: E = 200 + 0.35x Total revenue: R = 1.20x

OpenStudy (anonymous):

sorry im not good at math.

OpenStudy (mathstudent55):

The break even point is the number of cones you need to sell to cover all your expenses, but that gives you no profit.

OpenStudy (mathstudent55):

You're doing fine. Just keep paying attention.

OpenStudy (mathstudent55):

That point happens when the expenses equal the revenue.

OpenStudy (mathstudent55):

E = 200 + 0.35x R = 1.20x Now we set these two equations equal to each other and solve for x.

OpenStudy (anonymous):

ok

OpenStudy (mathstudent55):

200 + 0.35x = 1.20x Solve for x.

OpenStudy (anonymous):

kk

OpenStudy (mathstudent55):

200 + 0.35x = 1.20x Subtract 0.35x from both sides: 200 = 0.85x Now divide both sides by 0.85. x = 235.29 You need to sell 235.29 cones to break even. Since you can't sell 0.29 of a cone, you need to make 236 cones to break even.

OpenStudy (anonymous):

do i add 0.35+1.20x the divide it by 200?.... im sorry if thats wrong